OpenStudy (queenbb):

What is the equation of the graphed line? A graph with a line running through coordinates (-4, -6) and coordinates (2, 6)

OpenStudy (queenbb):

@RhondaSommer

OpenStudy (queenbb):

can you help me @OtherWorldly

OpenStudy (queenbb):

help plzzz

OpenStudy (otherworldly):

ok one sec let me see if i can help

OpenStudy (otherworldly):

i dont know how to get the equation sorry

OpenStudy (queenbb):

its ok

OpenStudy (queenbb):

@BlazeRyder

OpenStudy (blazeryder):

Let me look =)

OpenStudy (queenbb):

ok

OpenStudy (blazeryder):

Is there any particular equation? Point slope form, slope intercept form?

OpenStudy (queenbb):

its tells you the cordnites

RhondaSommer (rhondasommer):

\[\frac{12}8\\ \frac{3}2\\ y=\frac{3}2x+b\\ 6=\frac{3}2(2)+b\\ 6=\frac{6}2+b\\ 6=3+b\\ b=3\\ y=\frac{3}2x+3\]

RhondaSommer (rhondasommer):

there wa go :>

OpenStudy (queenbb):

thxs

RhondaSommer (rhondasommer):

np